f(x)=x+1 and g(x)=2x
find: f(g(x)) and g(f(x))

Answer by sdmmadam@yahoo.com(530)

f(x)=x+1----(1)and
g(x)=2x ----(2)
To find 1)f(g(x)) and 2)g(f(x))
1)f(g(x)) = f(2x)(using (2)
=f(t) where t = 2x ----(*)
=t+1 using (1)
=2x+1
2)g(f(x))
=g(x+1) using(1)
=g(u) where u = (x+1)----(**)
=2u using (2)
=2(x+1) using (**)
Answer: f(g(x))=2x+1
and g(f(x))=2(x+1)